# Beyond Aesthetics: Cultural Competence in Text-to-Image Models

### v5

@@ Appendix C Background on seed variation in T2I models
Appendix D Within-Culture Artifact Diversity
@@
Appendix E Proof of the scaling property.
@@
Appendix F Sensitivity of VS to prompt rephrasing and random seeds selection
@@
Appendix G Input Prompt Templates
@@
Appendix H Diversity Evaluation Pipeline
H.1 Prompting and Seeding
We utilize a simple prompting strategy tailored to our two research questions. For global geo-cultural diversity, prompts mention only the concept (e.g., "Image of monuments"), allowing us to measure the global inclination of model as well as the geo-cultural diversity of the generated artifacts. For evaluating within-culture diversity, prompts specify both the concept and the culture (e.g., "Image of a Nigerian dish"), to assess the cultural richness of a model’s representations within a specific cultural context. We use five prompt template variants for each run to account for potential variations in wording, to account for results dependence on prompt variance. Figure 7 shows the variance of VS across prompts templates.
For each prompt, we generate a batch of eight images using consecutive seed values. While our diversity metric is capable of handling larger batches, we limit our analysis to eight images, aligning with the typical output of image-generation APIs that present users with four or eight images per prompt. This process is repeated ten times, initiating at seed value 0, resulting in 80 unique seed values per prompt, spanning the range from 0 to 79. The mean diversity score is then calculated across all prompts rephrases and seed batches. Consequently, for each prompt, a total of 400 images are analyzed (5 prompts x 10 seed batches x 8 images/batch), and the scores are mean of 50 repetitions.
H.2 Mapping Generated Images to a Cultural Artifact from CUBE-CSpace
In this step, we map each generated image to a single cultural artifact within the prompt’s concept space. While not all text-to-image generated images perfectly represent real-world cultural entities, we aim to identify the most closely resembling artifact for each image. To ensure a clear mapping, we employ negative prompting during image generation to encourage the generation of images containing a single, dominant artifact. We further validate this using GPT-4-Turbo by prompting it to confirm whether the image indeed showcases a single prominent artifact.
For global geo-cultural diversity, where prompts focus on global concepts, we first use GPT-4-Turbo to verify that the generated image aligns with the concept in the prompt. Next, we utilize GPT-4-Turbo to identify the country most closely associated with the depicted artifact, emphasizing that we seek the country of prevalence or association, not necessarily origin. For within-culture diversity, where prompts specify a particular culture, we first use GPT-4-Turbo to confirm that the generated image aligns with the culture in the prompt. We leverage GPT-4-Turbo for these cultural faithfulness tasks, as it has been shown to be proficient in identifying cultural concepts (Cao et al., 2024). For both global and within culture diversity, we then retrieve the top 5 most similar images from a reference set of cultural artifact images for the identified country using the mSigLIP S400m model (Zhai et al., 2023) for image-image similarity. This model, trained on a global image dataset, is well-suited for our cultural analysis. The reference set consists of images from Google Images888 https://developers.google.com/custom-search/v1/overview for the artifacts in the concept space of the prompt. We acknowledge the limitations of Google Images as a proxy in the absence of a more authoritative global cultural artifact database. Finally, we employ GPT-4-Turbo and prompt it to classify the generated image as one of the five retrieved artifacts, mitigating potential noise from purely similarity-based retrieval. We human validate the entire mapping process on a small subset of images across cultures which results in $\sim$70 accuracy. We leave exploration of improved strategies to perform this mapping, potentially via more multicultural annotator models (VLMs), for future work.
H.3 Computing cultural diversity
To quantify the cultural diversity for a set of generated images, we employ the normalized qVS metric described in Section 5.1. This metric incorporates both the diversity of represented cultural artifacts and the quality of the generated images. For the latter, we leverage the HPS-v2 metric (Wu et al., 2023), a state-of-the-art metric for evaluating text-to-image generation based on human preferences. HPS-v2 captures crucial aspects of both image quality and faithfulness, effectively reflecting the faithfulness and aesthetic appeal of the generated images. While the HPS-v2 training data may not encompass the full spectrum of long-tail cultural artifacts considered in this work, it represents the most comprehensive and robust metric available for assessing human preferences in image generation, having been trained on an extensive dataset of 790,000 human preference ratings. In the absence of established datasets and evaluation models specifically tailored for cultural contexts, we adopt HPS-v2 as a proxy for overall generation quality. For the diversity part of the metric, we define different kernel functions that capture different aspects of geo-cultural diversity. We discuss more about the kernels in the Section 5.2.2.
H.4 Models Settings

### v6

@@ Appendix H Background on seed variation in T2I models
Appendix I Cultural Diversity Pipeline
We seek to analyze the global geo-cultural diversity of generated artifacts for under-specified prompts (Hutchinson et al., 2022) that simply mention the concept, such as "Image of traditional clothing.". These prompts serve as great test-beds to analyse model’s intrinsic cultural diversity. We further analyze within-culture diversity (What is the diversity of cultural artifacts produced by model for different cultures?) for Imagen 2 in Section J.
I.1 Prompting and Seeding
We employ a straightforward prompting strategy tailored to our two research questions on geo-cultural diversity. Figure 2 (bottom half) shows some example cultural diversity prompts. For evaluating global geo-cultural diversity, prompts consist only of the target concept (e.g., “Image of monuments”), enabling measurement of both the model’s global cultural inclination and the geo-cultural diversity in its generated artifacts. To assess within-culture diversity, prompts specify both the concept and the culture (e.g., “Image of a Nigerian dish”), allowing for analysis of the model’s cultural richness within specific cultural contexts.
To account for prompt wording and seed dependency, we use five distinct prompt templates and generate a batch of eight images per template with consecutive seed values, beginning with seed 0. This batch size aligns with typical outputs from image-generation APIs, which usually produce four or eight images per prompt. This process is repeated across ten seed batches per prompt, yielding a total of 80 unique seed values (0 to 79) for each prompt. The diversity metric, capable of processing larger batches, is computed over the 400 images generated per prompt (5 prompts x 10 seed batches x 8 images/batch). The resulting diversity scores represent the mean over 50 repetitions to ensure robustness. Figure 6 illustrates the sensitivity of diversity to prompt variation.
I.2 Computing Cultural Diversity
I.2.1 Mapping generated images to cultural artifacts
To compute the cultural diversity of the generated images, we first map each generated image to its closest resembling artifact within CUBE-CSpace. This mapping is essential to anchor our analysis in real-world cultural artifacts, even though we acknowledge that not all text-to-image generated images may perfectly represent actual cultural entities [^6] Generated images may deviate from real-world cultural representations.. Given that our prompts are designed to focus on broad global concepts, this approach allows us to associate each generated image $I$ with its corresponding continent $c$, country $r$, and artifact name $a$, such that $I\in\{c,r,a\}$.
For the mapping process, we employ an automated method that combines GPT-4-Turbo for verification with mSigLIP (Zhai et al., 2023)-based retrieval techniques. In cases where generated images might contain multiple artifacts, we use negative prompting to encourage the depiction of a single, dominant artifact. GPT-4-Turbo is then used to validate that each image contains a clearly identifiable primary artifact. For global geo-cultural diversity, where prompts describe global concepts, we use GPT-4-Turbo to confirm that the generated image aligns with the target concept. Following this, GPT-4-Turbo identifies the country most closely associated with the artifact, focusing on the country of prevalence or association rather than the origin. In the within-culture diversity case, where prompts specify both concept and culture, GPT-4-Turbo verifies that the generated image aligns with the specified culture in the prompt. This multi-step verification leverages GPT-4-Turbo’s proficiency in recognizing cultural concepts (Cao et al., 2024). For both global and within-culture diversity analyses, we retrieve the top five most similar images from a reference set of cultural artifact images associated with the identified country. This retrieval process leverages the mSigLIP S400m model (Zhai et al., 2023) for image-image similarity, which has been trained on a comprehensive global image dataset and is thus well-suited for this type of cultural analysis. The reference set comprises images sourced from Google Images [^7] https://developers.google.com/custom-search/v1/overview for artifacts within the prompt’s concept space. We recognize that Google Images, though extensive, may not represent the full range of global cultural artifacts. Finally, GPT-4-Turbo classifies each generated image by comparing it to the five retrieved reference artifacts, refining our results by reducing reliance on purely similarity-based retrieval. The entire mapping process is further validated by human reviewers on a small subset of images across diverse cultures, yielding an approximate accuracy of $\sim 70\%$. Exploration of improved mapping strategies, potentially using multicultural visual language models (VLMs) or diverse annotator models, is left as an avenue for future work.
I.2.2 Kernel Definition
With each generated image mapped to its closest cultural artifact, we compute the cultural diversity (CD) of the model’s output using the definition from Section 5.1. We define a general similarity kernel to analyze various aspects of geo-cultural diversity:
where $k_{1}(\cdot,\cdot)$, $k_{2}(\cdot,\cdot)$, and $k_{3}(\cdot,\cdot)$ are three distinct kernels representing different aspects of similarity, and $w_{1}$, $w_{2}$, $w_{3}$ assign weights to each. Specifically, $k_{1}(x_{i},x_{j})=1$ if $x_{i}$ and $x_{j}$ share the same continent, and 0 otherwise. Similarly, $k_{2}(x_{i},x_{j})=1$ if the items share the same country, and 0 otherwise. Lastly, $k_{3}(x_{i},x_{j})=1$ if the items represent the same artifact, irrespective of geographical origin, and 0 otherwise.
To demonstrate the flexibility of this kernel, we analyze cultural diversity using the following configurations:
  • Artifact-level diversity: $w_{1}=0$, $w_{2}=0$, $w_{3}=1$. This disregards geographical associations and measures diversity based solely on distinct artifacts.
  • Hierarchical geographical diversity: $w_{1}=\frac{1}{2}$, $w_{2}=\frac{1}{2}$, $w_{3}=0$. This captures a hierarchical notion of diversity, balancing both continent and country similarities equally without accounting for individual artifacts.
  • Uniformly weighted diversity: $w_{1}=\frac{1}{3}$, $w_{2}=\frac{1}{3}$, $w_{3}=\frac{1}{3}$. This provides equal weight to all three forms of similarity.
I.2.3 HPS-v2 to measure quality
To quantify the cultural diversity in a set of generated images, we employ the normalized qVS metric described in Section 5.1. This metric combines both the diversity of represented cultural artifacts and the quality of generated images. For the latter, we leverage the HPS-v2 metric (Wu et al., 2023), a state-of-the-art metric for evaluating text-to-image generation based on human preferences. HPS-v2 captures key aspects of image quality and faithfulness, effectively reflecting both the accuracy and aesthetic appeal of generated images. While HPS-v2’s training data may not fully encompass the long-tail cultural artifacts considered in this work, it remains the most comprehensive and robust metric available for assessing human preferences in image generation, having been trained on a dataset of 790,000 human preference ratings. In the absence of datasets and evaluation models specifically designed for cultural contexts, we adopt HPS-v2 as a proxy for overall generation quality. For the diversity component of the metric, we apply the different kernel functions defined above to capture distinct aspects of geo-cultural diversity. We provide additional details on the computation and application of these kernels in Section I.2.
I.2.4 Models Evaluated
Appendix J Within-Culture Artifact Diversity
@@ I.2.4 Models Evaluated
Appendix K Proof of the scaling property.
@@
Appendix L Sensitivity of VS to prompt rephrasing and random seeds selection
@@
Appendix M Input Prompt Templates